OpenStudy (anonymous):

The length of a rectangle plus its width is 24 cm. The area is 143 square cm. What are the length and width of the rectangle?

myininaya (myininaya):

l+w=24 lw=143=> w=143/l so we have l+143/l=24 l^2+143=24l l^2-24l+143=0

myininaya (myininaya):

(l-13)(l-14)=0 l=13,l=14

myininaya (myininaya):

so l=13,w=14 or l=14,w=13

OpenStudy (anonymous):

Thanks for the help :) !